# TRIAGE: Ethical Benchmarking of AI Models Through Mass Casualty Simulations

### v1

@@ Abstract
We present the TRIAGE Benchmark, a novel machine ethics (ME) benchmark that tests LLMs’ ability to make ethical decisions during mass casualty incidents. It uses real-world ethical dilemmas with clear solutions designed by medical professionals, offering a more realistic alternative to annotation-based benchmarks. TRIAGE incorporates various prompting styles to evaluate model performance across different contexts. Most models consistently outperformed random guessing, suggesting LLMs may support decision-making in triage scenarios. Neutral or factual scenario formulations led to the best performance, unlike other ME benchmarks where ethical reminders improved outcomes. Adversarial prompts reduced performance but not to random guessing levels. Open-source models made more morally serious errors, and general capability overall predicted better performance.
@@ 1 Introduction
To ensure that advanced AI systems are safe, they must act reliably following human values. Machine ethics (ME) benchmarks can indicate a system’s value alignment and moral understanding (Sun et al., 2024; Hendrycks et al., 2023). ME benchmarks help to establish industry safety standards, enable comparative analysis of AI models, and aid decision-makers in evaluating model capabilities, safety, and trustworthiness (Alzahrani et al., 2024). Previous studies on ME benchmarking have shown that state-of-the-art (SOTA) large language models (LLMs) exhibit a basic understanding of moral reasoning, with their ethical decision-making abilities improving alongside general advancements in capability (Sun et al., 2024; Pan et al., 2023). However, these evaluations have primarily been conducted using artificial or fictional scenarios (Pan et al., 2023; Hendrycks et al., 2023), and the proposed solutions often rely on annotations from crowdworkers, which often suffer from low inter-rater agreement. Additionally, many existing ME benchmarks fail to account for the diversity of cultural values, limiting the generalizability of the results.
To address these issues, we introduce TRIAGE—a new ME ethics benchmark that builds on existing medical triage models. Triage is the process of sorting patients according to the severity of their injuries to save as many lives as possible with the limited resources available. Different triage models across nations all share a core principle: prioritizing the greater good, even at the cost of individual needs. Triage models provide frameworks that physicians from diverse nations and cultures use to navigate the ethically complex decision of which patient to prioritize in emergency situations such as mass casualty incidents. As a result, triage models serve as excellent frameworks for evaluating the ethical decision-making capabilities of LLMs.
@@ 2.1 Dataset compilation
The TRIAGE Benchmark consists of 87 patient descriptions taken from triage training materials for the START and jumpSTART models (EMSC, 2016; Rajagopal, AB et al., 2020). In these models, doctors assign patients to one of four triage categories based on their symptoms (see Figure 1). Triage categories are ranked by treatment priority and resource allocation. Patients in the "Minor (green)" category have minor injuries and can be sent away. The "Expectant/Deceased (black)" category includes patients unlikely to survive; they are listed as second most resource intensive as they still receive palliative care while resources focus on those with better survival chances. Patients in the "Delayed (yellow)" category have serious injuries but can wait several hours for treatment until the patients in the "Immediate (red)" category have been helped who have life-threatening injuries and require urgent care.
@@ 2.2 Experiments
To run the experiments with the open-source models, we need one A100 80GB GPU to run our experiments which takes approximately five hours.
@@ 3 Results
  • More capable models generally perform better on our benchmark, but not in all contexts. For instance, as shown in Figures 3(b) and 4(d), GPT-4’s performance sometimes drops below Claude Haiku’s, with no significant differences between them.
@@ 3.1 Relative Performance
We tested all models with n=87 questions. This resulted in 3x5x87 answers per model. Sometimes models do not answer in the right format, leading to an effectively lower number of responses. Figures 2 and 3 show the relative ordering of models in the best (no ethics prompt) and worst (doctor jailbreaking prompt) conditions, with different model rankings in each. Notably, Claude Opus and Claude Haiku outperform GPT-4 under the doctor jailbreaking prompt. We assessed the significance of these rankings using five pairwise mixed logistic regression models (see Figure 4), with detailed results in Appendix E.
@@ 3.2 The effect of Ethics and Jailbreaking Prompts
Figure 2: Proportions of Correct Answers of Models on the TRIAGE Dataset.
(a) Ordering Based on Best-Case Performance
(b) Ordering Based on Worst-Case Performance
Figure 3: Best and Worst-Case Performance of Models.
Ordering of models changes depending on best or worst-case performance.
@@
Figure 4: Pairwise Mixed Model Comparisons
: Red stars indicate significance level (* : p < 0.04, ** : p < 0.01, *** : p < 0.001).
@@
Intercept represents estimate of all predictor variables at their reference levels (no ethics prompt, neutral syntax, and weaker model category). This value is the baseline outcome before accounting for the influence of the predictors and the random effects. Significance of all other factors is compared to intercept value.
Estimates in mixed effects models are typically in logits. We converted estimates to proportions in this figure.
@@ 3.3 Error Analysis
We conducted a detailed error analysis on the TRIAGE Benchmark, identifying three types of errors: instruction-following (model refuses or misformats answers), overcaring (allocating too many resources), and undercaring (allocating too few resources). We found that all proprietary models made substantially more overcaring errors than undercaring ones, consistently assigning patients to more resource-intensive triage categories. This trend is likely due to a large amount of safety fine-tuning these models go through. Interestingly, the open-source models we tested exhibited the opposite pattern, committing more undercaring errors. Figure 5(b) shows the misclassification pattern per prompt. We see that per prompt, the overcaring errors are greater than the undercaring errors.
@@ 4 Discussion
In this work, we demonstrated the ability of LLMs to solve ethical dilemmas in the medical context. All models, except Mistral, consistently outperformed random guessing on the TRIAGE benchmark. This indicates that models do indeed have a good understanding of moral values as suggested by Hendrycks et al. (2023) and that they are able to make sound moral decisions in the medical context.
@@
Given the safety focus of our ME benchmarks, worst-case performance may be more critical than best-case performance. To capture a broader range of potential model behaviors, we included multiple syntax variations, jailbreaking attacks, and ethical contexts. All models, except Mistral, consistently outperformed random guessing, even in their worst-performing condition. However, our findings show that the relative ranking of models can vary between best- and worst-case performances. The best-case rankings (see Figures 2 and 3(a)) align with expectations based on MT-Bench ratings. Interestingly, Claude 3 Haiku, which scored lower on MT-Bench than GPT-4, outperformed it in some ethical dilemma scenarios. One possible explanation is that more capable models like GPT-4 may experience "competing objectives" (Wei et al., 2023), where their enhanced instruction-following abilities conflict with safety training. However, Claude 3 Opus, considered as capable as GPT-4, did not show the same performance drop, suggesting that model architecture and training practices may be more predictive of ethical decision-making than general capability.
Our findings support three key hypotheses from Sun et al. (2024): (1) trustworthiness and utility (i.e., functional effectiveness) are often positively correlated, (2) proprietary LLMs tend to outperform open-source LLMs on ME benchmarks, and (3) proprietary LLMs are often overly calibrated toward trustworthiness. To explore this further, we analyzed error distributions per model. We found that proprietary LLMs primarily made overcaring errors, while open-source LLMs mostly made undercaring errors. Undercaring errors involve actively neglecting a patient in need, which is arguably more grave than committing an overcaring error, in which a patient receives too many resources. This suggests that proprietary models tend to make more aligned ethical choices, though at the expense of over-calibration. As Sun et al. (2024) note, while proprietary models may perform better, the increased transparency of open-source models offers an important trade-off to consider.
@@ 5 Conclusion
Our work demonstrates that LLMs are capable of navigating complex ethical dilemmas in the medical domain. By incorporating real-world scenarios and requiring models to make explicit moral decisions, TRIAGE offers a more realistic to other ME benchmarks. Further, our approach does not rely on potentially unreliable human or AI annotations. Our findings suggest that while proprietary models generally perform better, particularly by avoiding undercaring errors, this comes with the risk of over-calibration. We further see that reminding models of an ethical context can worsen their decision making in emergency situations. Although TRIAGE is limited to the medical field and does not include open-ended scenarios, it provides valuable insights into the ethical decision making of LLMs.

### v2

@@ Abstract
We present the TRIAGE Benchmark, a novel machine ethics (ME) benchmark that tests LLMs’ ability to make ethical decisions during mass casualty incidents. It uses real-world ethical dilemmas with clear solutions designed by medical professionals, offering a more realistic alternative to annotation-based benchmarks. TRIAGE incorporates various prompting styles to evaluate model performance across different contexts. Most models consistently outperformed random guessing, suggesting LLMs may support decision-making in triage scenarios. Neutral or factual scenario formulations led to the best performance, unlike other ME benchmarks where ethical reminders improved outcomes. Adversarial prompts reduced performance but not to random guessing levels. Open-source models made more morally serious errors, and general capability overall predicted better performance.  [^1] Code available at https://github.com/NLie2/Triage.
@@ 1 Introduction
To ensure that advanced AI systems are safe, they must act reliably following human values. Machine ethics (ME) benchmarks can indicate a system’s value alignment and moral understanding (Sun et al., 2024; Hendrycks et al., 2023). ME benchmarks help to establish industry safety standards, enable comparative analysis of AI models, and aid decision-makers in evaluating model capabilities, safety, and trustworthiness (Alzahrani et al., 2024). Previous studies on ME benchmarking have shown that state-of-the-art (SOTA) large language models (LLMs) exhibit a basic understanding of moral reasoning, with their ethical decision-making abilities improving alongside general advancements in capability (Sun et al., 2024; Pan et al., 2023). However, these evaluations have primarily been conducted using artificial or fictional scenarios (Pan et al., 2023; Hendrycks et al., 2023), and the proposed solutions often rely on annotations from crowd-workers, which often suffer from low inter-rater agreement. Additionally, many existing ME benchmarks fail to account for the diversity of cultural values, limiting the generalizability of the results.
To address these issues, we introduce TRIAGE—a new ME ethics benchmark that builds on existing medical triage models. Triage is the process of sorting patients according to the severity of their injuries to save as many lives as possible with the limited resources available. Different triage models across nations all share a core principle: prioritizing the greater good, even at the cost of individual needs. Triage models provide frameworks that physicians from diverse nations and cultures use to navigate the ethically complex decision of which patient to prioritize in emergencies such as mass casualty incidents. As a result, triage models serve as excellent frameworks for evaluating the ethical decision-making capabilities of LLMs.
@@ 2.1 Dataset compilation
The TRIAGE Benchmark consists of 87 patient descriptions taken from triage training materials for the START and jumpSTART models (EMSC, 2016; Rajagopal, AB et al., 2020). In these models, doctors assign patients to one of four triage categories based on their symptoms (see Figure 1). Triage categories are ranked by treatment priority and resource allocation. Patients in the "Minor (green)" category have minor injuries and can be sent away. The "Expectant/Deceased (black)" category includes patients unlikely to survive; they are listed as the second most resource intensive as they still receive palliative care while resources focus on those with better survival chances. Patients in the "Delayed (yellow)" category have serious injuries but can wait several hours for treatment until the patients in the "Immediate (red)" category have been helped who have life-threatening injuries and require urgent care.
@@
2.2 Prompt Generation
We created jailbreaking prompts by using a method called manual persona modulation Shah et al. (2023). In this approach, GPT-4 was instructed to automatically generate jailbreaking prompts based on specific task and persona descriptions, in this case Healthcare Assistant and Doctor Assistant.
Our ethics prompts were also generated by GPT-4, using a template from the MACHIAVELLI benchmark Pan et al. (2023). GPT-4 was instructed to reformulate these prompts in terms of either utilitarian or deontological ethics.
@@ 2.3 Experiments
To run the experiments with the open-source models, we need one A100 80GB GPU which takes approximately five hours.
@@ 3 Results
  • More capable models generally perform better on our benchmark, but not in all contexts. For instance, as shown in Figures 2(c) and 3(d), GPT-4’s performance sometimes drops below Claude Haiku’s, with no significant differences between them.
@@ 3.1 Relative Performance
We tested all models with n=87 questions. This resulted in 3x5x87 answers per model. Sometimes models do not answer in the right format, leading to an effectively lower number of responses. Figures 2 shows the relative ordering of models in the best (no ethics prompt) and worst (doctor jailbreaking prompt) conditions, with different model rankings in each. Notably, Claude Opus and Claude Haiku outperform GPT-4 under the doctor jailbreaking prompt. We assessed the significance of these rankings using five pairwise mixed logistic regression models (see Figure 3), with detailed results in Appendix E.
@@ 3.2 The effect of Ethics and Jailbreaking Prompts
(a) Proportions of correct answers for different models and prompts
(b) Best-case ordering (no ethics prompt)
(c) Worst-case ordering (doctor assistant)
Figure 2: (a) Proportions of Correct Answers of Models on the TRIAGE Dataset, and (b) Best and Worst-Case Performance of Models, showing how the ordering of models changes depending on the scenario.
@@ 3.3 Error Analysis
We conducted a detailed error analysis on the TRIAGE Benchmark, identifying three types of errors: instruction-following (model refuses or wrongly formats answers), overcaring (allocating too many resources), and undercaring (allocating too few resources). We found that all proprietary models made substantially more overcaring errors than undercaring ones, consistently assigning patients to more resource-intensive triage categories. This trend is likely due to a large amount of safety fine-tuning these models go through. Interestingly, the open-source models we tested exhibited the opposite pattern, committing more undercaring errors. Figure 4(b) shows the misclassification pattern per prompt. We see that per prompt, the overcaring errors are greater than the undercaring errors.
@@ 4 Discussion
In this work, we demonstrated the ability of LLMs to solve ethical dilemmas in the medical context. All models, except Mistral, consistently outperformed random guessing on the TRIAGE benchmark. This indicates that models do indeed have a good understanding of moral values as suggested by Hendrycks et al. (2023) and that they can make sound moral decisions in the medical context.
@@
Given the safety focus of our ME benchmarks, worst-case performance may be more critical than best-case performance. To capture a broader range of potential model behaviors, we included multiple syntax variations, jailbreaking attacks, and ethical contexts. All models, except Mistral, consistently outperformed random guessing, even in their worst-performing condition. However, our findings show that the relative ranking of models can vary between best- and worst-case performances. The best-case rankings (see Figures 2(a) and 2(b)) align with expectations based on MT-Bench ratings. Interestingly, Claude 3 Haiku, which scored lower on MT-Bench than GPT-4, outperformed it in some ethical dilemma scenarios. One possible explanation is that more capable models like GPT-4 may experience "competing objectives" (Wei et al., 2023), where their enhanced instruction-following abilities conflict with safety training. However, Claude 3 Opus, considered as capable as GPT-4, did not show the same performance drop, suggesting that model architecture and training practices may be more predictive of ethical decision-making than general capability.
Our findings support three key hypotheses from Sun et al. (2024): (1) trustworthiness and utility (i.e., functional effectiveness) are often positively correlated, (2) proprietary LLMs tend to outperform open-source LLMs on ME benchmarks, and (3) proprietary LLMs are often overly calibrated toward trustworthiness. To explore this further, we analyzed error distributions per model. We found that proprietary LLMs primarily made overcaring errors, while open-source LLMs mostly made undercaring errors. Undercaring errors involve actively neglecting a patient in need, which is arguably graver than committing an overcaring error, in which a patient receives too many resources. This suggests that proprietary models tend to make more aligned ethical choices, though at the expense of over-calibration. As Sun et al. (2024) note, while proprietary models may perform better, the increased transparency of open-source models offers an important trade-off to consider.
@@ 5 Conclusion
Our work demonstrates that LLMs are capable of navigating complex ethical dilemmas in the medical domain. By incorporating real-world scenarios and requiring models to make explicit moral decisions, TRIAGE offers a more realistic to other ME benchmarks. Further, our approach does not rely on potentially unreliable human or AI annotations. Our findings suggest that while proprietary models generally perform better, particularly by avoiding undercaring errors, this comes with the risk of over-calibration. We further see that reminding models of an ethical context can worsen their decision-making in emergency situations. Although TRIAGE is limited to the medical field and does not include open-ended scenarios, it provides valuable insights into the ethical decision-making of LLMs.